# FRRI: a novel algorithm for fuzzy-rough rule induction

### v1

@@ Abstract
Interpretability is the next frontier in machine learning research. In the search for white box models — as opposed to black box models, like random forests or neural networks — rule induction algorithms are a logical and promising option, since the rules can easily be understood by humans. Fuzzy and rough set theory have been successfully applied to this archetype, almost always separately. As both approaches to rule induction involve granular computing based on the concept of equivalence classes, it is natural to combine them. The QuickRulesJensenCornelis2009 algorithm was a first attempt at using fuzzy rough set theory for rule induction. It is based on QuickReduct, a greedy algorithm for building decision reducts. QuickRules already showed an improvement over other rule induction methods. However, to evaluate the full potential of a fuzzy rough rule induction algorithm, one needs to start from the foundations. In this paper, we introduce a novel rule induction algorithm called Fuzzy Rough Rule Induction (FRRI). We provide background and explain the workings of our algorithm. Furthermore, we perform a computational experiment to evaluate the performance of our algorithm and compare it to other state-of-the-art rule induction approaches. We find that our algorithm is more accurate while creating small rulesets consisting of relatively short rules. We end the paper by outlining some directions for future work.
@@ 1 Introduction
Many fuzzy rule induction algorithms have been established, mostly for deriving a concise set of rules comprehensible by humans for tasks like classification and prediction. These include, for example, fuzzy association rule mining Dong2005 ; 1593646 , first-order fuzzy rule generation PradeRIchardSerr2003 ; DROBICS2003131 , and linguistic semantics-preserving modeling MarnBlzquez2002FromAT ; QIN2008435 . However, the efficacy of most of the existing approaches to fuzzy rule induction is reduced as the data dimensionality increases. Some methods manage to avoid this, for example standard covering algorithms for rule induction (e.g., RIPPER COHEN1995115 and its fuzzy counterpart, FURIA FURIA ) that learn rules in an incremental way, with each rule in turn constructed by adding maximally informative features one by one.
QuickRules JensenCornelis2009 was developed to combat this problem by directly integrating the rule induction process in a greedy feature selection algorithm based on fuzzy rough set theory fuzzyroughsets ; jensen2008computational . Despite promising initial results reported in JensenCornelis2009 , the greedy nature of this algorithm does not allow it to outperform state-of-the-art algorithms w.r.t. balanced accuracy, ruleset size and rule length, as we will also demonstrate in the experimental section of this paper.
@@ 2.1 Fuzzy sets
In this paper, $U$ will indicate a finite, non-empty set, called the universe. A fuzzy set fuzzysets $A$ in $U$ is a function $A\colon U\to[0,1]$. For $x$ in $U$, $A(x)$ is called the membership degree of $x$ in $A$. The set of all fuzzy sets of $U$ is denoted as $\mathscr{F}(U)$. Let $A,B\in\mathscr{F}(U)$ be two fuzzy sets. $A$ is a fuzzy subset of $B$ if
@@
A (binary) fuzzy relation on a universe $U$ is simply a fuzzy subset of $U\times U$. A fuzzy relation $R$ is reflexive if $R(x,x)=1$ for all elements $x\in U$. It is symmetric if for every pair of elements $x$, $y$ in the universe, $R(x,y)=R(y,x)$. For a given t-norm $\operatorname{\mathcal{T}}$, $R$ is $\operatorname{\mathcal{T}}$-transitive if for every triplet of elements $x,y,z$, $\operatorname{\mathcal{T}}(R(x,y),R(y,z))\leq R(x,z)$. A reflexive and $\operatorname{\mathcal{T}}$-transitive relation is called a $\operatorname{\mathcal{T}}$-preorder relation. A fuzzy tolerance relation is a reflexive and symmetric fuzzy relation. Given an element $x\in U$, its fuzzy tolerance class w.r.t. to a tolerance relation $R$ is the fuzzy set $Rx(y)=R(x,y)$, for any $y\in U$.
@@ 2.2 Fuzzy rough sets
In classical rough set theory roughsets , we examine a universe divided into equivalence classes by an equivalence relation. We can then approximate a concept (i.e., a subset of $U$) using two crisp sets: the lower and the upper approximation. The lower approximation contains all elements which are certainly part of the concept. It is defined as the union of the equivalence classes which are subsets of the concept. The upper approximation, which contains all elements which might possibly be a part of the concept, is defined as the union of all equivalence classes which have a non-empty intersection with the concept in question. In fuzzy rough sets fuzzyroughsets , we replace the crisp equivalence relation with a binary fuzzy relation $R$ on $U$ and, in turn, the upper and lower approximations also become fuzzy sets. Let $A$ be a fuzzy set in $U$. Furthermore, let $\operatorname{\mathcal{T}}$ be a t-norm and $\operatorname{\mathcal{I}}$ be an implicator. We define the fuzzy rough set (FRS) for the fuzzy set $A$ w.r.t. the fuzzy relation $R$ as the pair $(\underline{A}_{R}^{\operatorname{\mathcal{I}}},\overline{A}_{R}^{\operatorname{\mathcal{T}}})$, with, for $x$ in $U$:
@@
    $\underline{A}_{R}^{\operatorname{\mathcal{I}}}(x)=\min_{y\in U}\operatorname{\mathcal{I}}(R(x,y),A(y))$ (1)
@@
    $\overline{A}_{R}^{\operatorname{\mathcal{T}}}(x)=\max_{y\in U}\operatorname{\mathcal{T}}(R(x,y),A(y))$ (2)
@@ 2.3 Information and decision systems
An information system roughsets is a pair $(U,\operatorname{\mathcal{A}})$, with $U$ a finite, non-empty universe of objects, and $\operatorname{\mathcal{A}}$ a finite, non-empty set of attributes describing these objects. Each attribute $a\in\operatorname{\mathcal{A}}$ has an associated domain of possible values $V_{a}$ and an associated function $a\colon U\to V_{a}$ which maps an object $x\in U$ to its value $a(x)$ for attribute $a$. An attribute is numerical if its value set is a closed real interval. In the remainder of this paper, we will only work in normalised information systems with numerical condition attributes, where all attributes are scaled such that they have unit range, i.e., such that $V_{a}=[0,1]$ for all attributes $a\in\operatorname{\mathcal{A}}$. This scaling step is important because the original range of a condition attribute often has no bearing on the true importance of the feature as a predictor for the class. To this aim, we will preprocess information systems by performing minimum-maximum normalisation, i.e., we replace each original attribute function $a\in\operatorname{\mathcal{A}}$ with the following alternate attribute function $a^{\prime}$:
@@
$R_{B}(x,y)=\operatorname{\mathcal{T}}(\underbrace{R_{a}(x,y)}_{a\in B})$ (5)
for each $x$ and $y$ in $U$ and for a given t-norm $\operatorname{\mathcal{T}}$, where $R_{a}$ is the relation used for comparing the values of attribute $a$. In general, $R_{B}$ is a reflexive fuzzy relation.
A decision system $(U,\operatorname{\mathcal{A}}\cup\{d\})$ with a single decision attribute $d$ is an information system which makes a distinction between the condition attributes $\operatorname{\mathcal{A}}$ and the decision attribute $d$, which is not an element of $\operatorname{\mathcal{A}}$. In this paper, the decision attribute is assumed to be categorical, which means that it has a finite, unordered value domain. We will call the values in $V_{d}$, decision classes. For a given object $u$ in $U$, we will identify its decision class $d(u)$ with the set of all objects that have the same value for the decision attribute. Given a subset of attributes $B\subseteq\operatorname{\mathcal{A}}$, the fuzzy $B$-positive region is a fuzzy set in the universe $U$ that contains each object $u$ to the extent that all objects which are approximately similar to $u$ w.r.t. the attributes in $B$, have the same decision value $d$ CORNELIS2010209 :
@@ 2.4 Rule induction
In the following sections, we will give a short description of some fuzzy and rough rule induction algorithms. RIPPER COHEN1995115 , and its fuzzified version FURIA FURIA , are state-of-the-art (fuzzy) rule induction algorithms that work in two phases. In the first phase, an initial ruleset is developed using a modified version of Incremental Reduced Error Pruning (IREP*) furnkranz1994incremental . Afterwards, in the second phase, an optimization is applied, which calculates possible alternatives for each rule, and selects the best one. To round out the optimization step, IREP* is once again used to fill in any possible gaps left by selecting alternate rules, this optimization step can be repeated multiple times.
@@
Among rough set based rule induction algorithms, MODLEM Stefanowski1998OnRS is generally considered to be the standard option due to its excellent performance. This algorithm was inspired by LEM2 threediscretmodlem . By contrast to the former, it has no need for a discretisation step, selecting the best cut-off values for numerical attributes during the rule generation. MODLEM follows a heuristic strategy for creating an initial rule by choosing sequentially the “best” elementary conditions according to some heuristic criteria. Learning examples that match this rule are removed from consideration. The process is repeated iteratively as long as some learning examples remain uncovered. The resulting set of rules covers all learning examples.
@@
There have been very few attempts at developing rule induction algorithms based on fuzzy rough set theory. A prominent example is QuickRulesJensenCornelis2009 , a greedy fuzzy rough algorithm based on the feature selection algorithm QuickReduct fuzzyroughsets ; jensen2008computational . It greedily selects the most discerning attribute according to its contribution to the positive region. During these calculations, candidate rules are checked and possibly added to a growing ruleset. Recently, the use of the Hamacher t-norm in the context of QuickRules has been explored in hadamacher .
@@
Other work has largely focused on using classical rough set theory to generate fuzzy rulesets Hsiesh2008 ; ShenChouchoulas02 , but mainly ignores the direct use of fuzzy-rough concepts. The induction of gradual decision rules, based on fuzzy-rough hybridization, has been presented in GRECO2006179 . For this approach, new definitions of fuzzy lower and upper approximations were constructed that avoid the use of fuzzy logical connectives altogether. In this approach, decision rules are induced from lower and upper approximations defined for positive and negative relationships between credibility of premises and conclusions. Only the ordinal properties of fuzzy membership degrees are used. Another fuzzy-rough approach to fuzzy rule induction was presented in WangTsangZhaoChen2007 , where fuzzy reducts are employed to generate rules from data. This method also employs a fuzzy-rough feature selection preprocessing step. In JensenShenbook2008 , a fuzzy decision tree algorithm was proposed, based on fuzzy ID3, that incorporates the fuzzy-rough dependency function as a splitting criterion. A fuzzy-rough rule induction method was also proposed in Hong2006 for generating certain and possible rulesets from hierarchical data. This differs from our approach, in which we will only generate a single ruleset, as we will explain in the next section.
@@ 3.1 Rule format
We also denote the fuzzy set $M_{r}$ as the matching set of $r$. The relations corresponding to each condition type are, for $x,y\in[0,1]$:
$R_{\textsc{similar}}(x,y)$ $=R_{i}(x,y)$
$R_{\textsc{dominant}}(x,y)$ $=R_{d}(x,y)$
$R_{\textsc{dominated}}(x,y)$ $=R_{d}(y,x)$
$R_{\textsc{unused}}(x,y)$ $=1$
We can see that dominant and dominated conditions are indeed more general than similar conditions, since, for the same $x\in[0,1]$, it holds that, for any $y\in[0,1]$,
$R_{\textsc{similar}}(x,y)\leq R_{\textsc{dominant}}(x,y)\text{ and }R_{\textsc{similar}}(x,y)\leq R_{\textsc{dominated}}(x,y)$
@@
Throughout this section, we will use the decision system presented in Table 1 to illustrate the operation of our algorithm. It is a sample of the Pima Indians Diabetes dataset taken from the KEEL dataset repository keeldatarepo , which we have normalised using min-max normalisation, such that all attributes have unit range. We consider a universe with seven objects, eight numerical condition attributes and one binary decision attribute. We consider an example rule $r$ derived from object $u_{1}$ from our example decision system:
@@ 3.3 Rule selection
Finally, we can formulate the optimization problem as follows:
@@
or in other words: minimize the number of selected rules, while still assuring that every object in the universe is covered by at least one of the selected rules. This integer programming problem can be solved with classical techniques, such as those found in the gurobipy package for Python gurobi .
@@ 4.1 Methods and evaluation measures
Table 3: Details of the benchmark datasets used for these experiments: abbreviated name, the amount of features of each dataset ($|\operatorname{\mathcal{A}}|$), the number of instances ($|U|$) and classes ($|V_{d}|$), and the imbalance ration IR
We perform a computational experiment to evaluate the performance of our novel FRRI algorithm on 18 benchmark numerical datasets from the KEEL dataset repository [^1] Some of these datasets usually have categorical condition features. We opted to remove these features from the dataset before using them in our experiments. keeldatarepo . The details of these datasets can be found in Table 3, where we list the amount of features of each dataset ($|A|$), as well as the number of instances ($|U|$) and classes ($|V_{d}|$). Finally, the table also contains the imbalance ratio (IR) of each dataset, which is calculated as the size of the largest class divided by the size of the smallest class. We also list an abbreviated name for each dataset.
We implement our algorithm in Python and also compare it with QuickRules JensenCornelis2009 , MODLEM Stefanowski1998OnRS , FURIA FURIA and RIPPER COHEN1995115 . We use our own Python implementation of QuickRules, and we use the implementation found in WEKAWEKA for the other algorithms. We also use the default parameters for all algorithms.
For each algorithm, we perform ten-fold cross-validation on each dataset, using the same folds for each algorithm, after which we calculate the average balanced accuracy of the algorithm on the ten training folds. The balanced accuracy is the arithmetic mean of the recall on each class. Moreover, we also calculate the average number of rules and the average rule length over each ruleset and over each of the ten folds of each dataset. To determine the statistical significance of our results, we apply a two-stage procedure, which starts with a Friedman test multiple-comp-friedman . If the result of that test is significant, we continue with a Conover post-hoc pairwise comparison procedure multiple-comp-conover ; conover_test_big_citations to determine the location of the significant differences. Moreover, we also use the Wilcoxon test pairwise-comp-wilcoxon to detect differences between two methods.

### v2

@@ Abstract
Interpretability is the next frontier in machine learning research. In the search for white box models — as opposed to black box models, like random forests or neural networks — rule induction algorithms are a logical and promising option, since the rules can easily be understood by humans. Fuzzy and rough set theory have been successfully applied to this archetype, almost always separately. As both approaches offer different ways to deal with imprecise and uncertain information, often with the use of an indiscernibility relation, it is natural to combine them. The QuickRules [20] algorithm was a first attempt at using fuzzy rough set theory for rule induction. It is based on QuickReduct, a greedy algorithm for building decision superreducts. QuickRules already showed an improvement over other rule induction methods. However, to evaluate the full potential of a fuzzy rough rule induction algorithm, one needs to start from the foundations. Accordingly, the novel rule induction algorithm, Fuzzy Rough Rule Induction (FRRI), we introduce in this paper, uses an approach that has not yet been utilised in this setting. We provide background and explain the workings of our algorithm. Furthermore, we perform a computational experiment to evaluate the performance of our algorithm and compare it to other state-of-the-art rule induction approaches. We find that our algorithm is more accurate while creating small rulesets consisting of relatively short rules.
@@ 1 Introduction
Many fuzzy rule induction algorithms have been established, mostly for deriving a concise set of rules comprehensible by humans for tasks like classification and prediction. These include, for example, fuzzy association rule mining [3, 31], first-order fuzzy rule generation [9, 25], and linguistic semantics-preserving modeling [22, 26]. However, the efficacy of most of the existing approaches to fuzzy rule induction is reduced as the data dimensionality increases. Some methods manage to avoid this, for example standard covering algorithms for rule induction (e.g., RIPPER [4] and its fuzzy counterpart, FURIA [19]) that learn rules in an incremental way, with each rule in turn constructed by adding maximally informative features one by one.
QuickRules [20] was developed to combat this problem by directly integrating the rule induction process in a greedy feature selection algorithm based on fuzzy rough set theory [7, 10, 21]. Despite promising initial results reported in [20], the greedy nature of this algorithm does not allow it to outperform state-of-the-art algorithms w.r.t. balanced accuracy, ruleset size and rule length, as we will also demonstrate in the experimental section of this paper.
@@ 2.1 Fuzzy sets
In this paper, $U$ will indicate a finite, non-empty set, called the universe. A fuzzy set [32] $A$ in $U$ is a function $A\colon U\to[0,1]$. For $u$ in $U$, $A(u)$ is called the membership degree of $u$ in $A$. The set of all fuzzy sets of $U$ is denoted as $\mathscr{F}(U)$. Let $A,B\in\mathscr{F}(U)$ be two fuzzy sets. $A$ is a fuzzy subset of $B$ if
@@
A (binary) fuzzy relation on a universe $U$ is simply a fuzzy subset of $U\times U$. A fuzzy relation $R$ is reflexive if $R(u,u)=1$ for all elements $u\in U$. It is symmetric if for every pair of elements $u$, $v$ in the universe, $R(u,v)=R(v,u)$. For a given t-norm $\operatorname{\mathcal{T}}$, $R$ is $\operatorname{\mathcal{T}}$-transitive if for every triplet of elements $u,v,w$, $\operatorname{\mathcal{T}}(R(u,v),R(v,w))\leq R(u,w)$. A reflexive and $\operatorname{\mathcal{T}}$-transitive relation is called a $\operatorname{\mathcal{T}}$-preorder relation. A fuzzy tolerance relation is a reflexive and symmetric fuzzy relation. Given an element $u\in U$, its fuzzy tolerance class w.r.t. to a tolerance relation $R$ is the fuzzy set $Ru(y)=R(u,v)$, for any $v\in U$.
@@ 2.2 Fuzzy rough sets
In classical rough set theory [24], we examine a universe divided into equivalence classes by an equivalence relation. We can then approximate a concept (i.e., a subset of $U$) using two crisp sets: the lower and the upper approximation. The lower approximation contains all elements which are certainly part of the concept. It is defined as the union of the equivalence classes which are subsets of the concept. The upper approximation, which contains all elements which might possibly be a part of the concept, is defined as the union of all equivalence classes which have a non-empty intersection with the concept in question. In fuzzy rough sets [10], we replace the crisp equivalence relation with a binary fuzzy relation $R$ on $U$ and, in turn, the upper and lower approximations also become fuzzy sets. Let $A$ be a fuzzy set in $U$. Furthermore, let $\operatorname{\mathcal{T}}$ be a t-norm and $\operatorname{\mathcal{I}}$ be an implicator. We define the fuzzy rough set (FRS) for the fuzzy set $A$ w.r.t. the fuzzy relation $R$ as the pair $(\underline{A}_{R}^{\operatorname{\mathcal{I}}},\overline{A}_{R}^{\operatorname{\mathcal{T}}})$, with, for $u$ in $U$:
@@
    $\underline{A}_{R}^{\operatorname{\mathcal{I}}}(u)=\min_{v\in U}\operatorname{\mathcal{I}}(R(u,v),A(v))$ (1)
@@
    $\overline{A}_{R}^{\operatorname{\mathcal{T}}}(u)=\max_{v\in U}\operatorname{\mathcal{T}}(R(u,v),A(v))$ (2)
@@ 2.3 Information and decision systems
An information system [24] is a pair $(U,\operatorname{\mathcal{A}})$, with $U$ a finite, non-empty universe of objects, and $\operatorname{\mathcal{A}}$ a finite, non-empty set of attributes describing these objects. Each attribute $a\in\operatorname{\mathcal{A}}$ has an associated domain of possible values $V_{a}$ and an associated function $a\colon U\to V_{a}$ which maps an object $u\in U$ to its value $a(u)$ for attribute $a$. An attribute is numerical if its value set is a closed real interval. In the remainder of this paper, we will only work in normalised information systems with numerical condition attributes, where all attributes are scaled such that they have unit range, i.e., such that $V_{a}=[0,1]$ for all attributes $a\in\operatorname{\mathcal{A}}$. This scaling step is important because the original range of a condition attribute often has no bearing on the true importance of the feature as a predictor for the class. To this aim, we will preprocess information systems by performing minimum-maximum normalisation, i.e., we replace each original attribute function $a\in\operatorname{\mathcal{A}}$ with the following alternate attribute function $a^{\prime}$:
@@
$R_{B}(u,v)=\operatorname{\mathcal{T}}(\underbrace{R_{a}(u,v)}_{a\in B})$ (5)
for each $u$ and $v$ in $U$ and for a given t-norm $\operatorname{\mathcal{T}}$, where $R_{a}$ is the relation used for comparing the values of attribute $a$. In general, $R_{B}$ is a reflexive fuzzy relation.
A decision system $(U,\operatorname{\mathcal{A}}\cup\{d\})$ with a single decision attribute $d$ is an information system which makes a distinction between the condition attributes $\operatorname{\mathcal{A}}$ and the decision attribute $d$, which is not an element of $\operatorname{\mathcal{A}}$. In this paper, the decision attribute is assumed to be categorical, which means that it has a finite, unordered value domain. We will call the values in $V_{d}$, decision classes. For a given object $u$ in $U$, we will identify its decision class $d(u)$ with the set of all objects that have the same value for the decision attribute. Given a subset of attributes $B\subseteq\operatorname{\mathcal{A}}$, the fuzzy $B$-positive region is a fuzzy set in the universe $U$ that contains each object $u$ to the extent that all objects which are approximately similar to $u$ w.r.t. the attributes in $B$, have the same decision value $d$ [8]:
@@ 2.4 Rule induction
In the following sections, we will give a short description of some fuzzy and rough rule induction algorithms. RIPPER [4], and its fuzzified version FURIA [19], are state-of-the-art (fuzzy) rule induction algorithms that work in two phases. In the first phase, an initial ruleset is developed using a modified version of Incremental Reduced Error Pruning (IREP*) [13]. Afterwards, in the second phase, an optimisation is applied, which calculates possible alternatives for each rule, and selects the best one. To round out the optimisation step, IREP* is once again used to fill in any possible gaps left by selecting alternate rules, this optimisation step can be repeated multiple times.
@@
Rough set theory is also widely applied to rule induction. An overview of applications of rough sets can be found here [2]. Among rough set based rule induction algorithms, MODLEM [28] is generally considered to be the standard option due to its excellent performance. This algorithm was inspired by LEM2 [15]. By contrast to the former, it has no need for a discretisation step, selecting the best cut-off values for numerical attributes during the rule generation. MODLEM follows a heuristic strategy for creating an initial rule by choosing sequentially the “best” elementary conditions according to some heuristic criteria. Learning examples that match this rule are removed from consideration. The process is repeated iteratively as long as some learning examples remain uncovered. The resulting set of rules covers all learning examples.
@@
There have been very few attempts at developing rule induction algorithms based on fuzzy rough set theory. A prominent example is QuickRules [20], a greedy fuzzy rough algorithm based on the feature selection algorithm QuickReduct [10, 21]. It greedily selects the most discerning attribute according to its contribution to the positive region. During these calculations, candidate rules are checked and possibly added to a growing ruleset. Recently, the use of the Hamacher t-norm in the context of QuickRules has been explored in [23].
@@
Other work has largely focused on using classical rough set theory to generate fuzzy rulesets [18, 27], but mainly ignores the direct use of fuzzy-rough concepts. The induction of gradual decision rules, based on fuzzy-rough hybridisation, has been presented in [14]. For this approach, new definitions of fuzzy lower and upper approximations were constructed that avoid the use of fuzzy logical connectives altogether. In this approach, decision rules are induced from lower and upper approximations defined for positive and negative relationships between credibility of premises and conclusions. Only the ordinal properties of fuzzy membership degrees are used. Another fuzzy-rough approach to fuzzy rule induction was presented in [29], where fuzzy reducts are employed to generate rules from data. This method also employs a fuzzy-rough feature selection preprocessing step. In [21], a fuzzy decision tree algorithm was proposed, based on fuzzy ID3, that incorporates the fuzzy-rough dependency function as a splitting criterion. A fuzzy-rough rule induction method was also proposed in [17] for generating certain and possible rulesets from hierarchical data. This differs from our approach, in which we will only generate a single ruleset, as we will explain in the next section.
@@ 3.1 Rule format
We also denote the fuzzy set $M_{r}$ as the matching set of $r$. The relations corresponding to each condition type are, for $u,v\in[0,1]$:
$R_{\textsc{similar}}(u,v)$ $=R_{i}(u,v)$
$R_{\textsc{dominant}}(u,v)$ $=R_{d}(u,v)$
$R_{\textsc{dominated}}(u,v)$ $=R_{d}(v,u)$
$R_{\textsc{unused}}(u,v)$ $=1$
We can see that dominant and dominated conditions are indeed more general than similar conditions, since, for the same $u\in[0,1]$, it holds that, for any $v\in[0,1]$,
$R_{\textsc{similar}}(u,v)\leq R_{\textsc{dominant}}(u,v)\text{ and }R_{\textsc{similar}}(u,v)\leq R_{\textsc{dominated}}(u,v)$
@@
Throughout this section, we will use the decision system presented in Table 1 to illustrate the operation of our algorithm. It is a sample of the Pima Indians Diabetes dataset taken from the KEEL dataset repository [1], which we have normalised using min-max normalisation, such that all attributes have unit range. We consider a universe with seven objects, eight numerical condition attributes and one binary decision attribute. We consider an example rule $r$ derived from object $u_{1}$ from our example decision system:
@@ 3.3 Rule selection
Finally, we can formulate the optimisation problem as follows:
@@
or in other words: minimize the number of selected rules, while still assuring that every object in the universe is covered by at least one of the selected rules. This integer programming problem can be solved with classical techniques, such as branch-and-bound algorithms.
@@ 4.1 Methods and evaluation measures
Table 3: Details of the benchmark datasets used for these experiments: abbreviated name, the amount of features of each dataset ($|\operatorname{\mathcal{A}}|$), the number of instances ($|U|$) and classes ($|V_{d}|$), and the imbalance ratio IR
We perform a computational experiment to evaluate the performance of our novel FRRI algorithm on 18 benchmark numerical datasets from the KEEL dataset repository [^1] Some of these datasets usually have categorical condition features. We opted to remove these features from the dataset before using them in our experiments. [1]. The details of these datasets can be found in Table 3, where we list the amount of features of each dataset ($|A|$), as well as the number of instances ($|U|$) and classes ($|V_{d}|$). Finally, the table also contains the imbalance ratio (IR) of each dataset, which is calculated as the size of the largest class divided by the size of the smallest class. We also list an abbreviated name for each dataset.
We implement our algorithm in Python, using the gurobipy package for Python [16] to solve the optimisation problem in the rule selection step. We chose Gurobi because it is a mature, well tested product with good support, free access for academic use, with a strong community and an excellent Python implementation. However, any other exact optimizer (including open-source alternatives) could be used without fundamentally changing the FRRI algorithm and the obtained experimental results, which were obtained after repeated runs.
We compare FRRI with QuickRules [20], MODLEM [28], FURIA [19] and RIPPER [4]. We use the implementation found in WEKA [11] for the latter algorithms and we use our own Python implementation of QuickRules, which was verified against the one in WEKA. We also use the default parameters for all algorithms.
For each algorithm, we perform ten-fold cross-validation on each dataset, using the same folds for each algorithm, after which we calculate the average balanced accuracy of the algorithm on the ten training folds. The balanced accuracy is the arithmetic mean of the recall on each class. Moreover, we also calculate the average number of rules and the average rule length over each ruleset and over each of the ten folds of each dataset. To determine the statistical significance of our results, we apply a two-stage procedure, which starts with a Friedman test [12]. If the result of that test is significant, we continue with a Conover post-hoc pairwise comparison procedure [5, 6] to determine the location of the significant differences. Moreover, we also use the Wilcoxon test [30] to detect differences between two methods.